OpenStudy (anonymous):

If the Federal Reserve decreases the reserve rate from 5% to 2.5%, how does this affect the amount of money that would result because of fractional-reserve banking from an initial deposit into a bank of $35,000? A. It incraeses the amount by 1,400,000 B.It decreases the amount by 1,400,000 C.It decreases the amount by 700,000 D. It increases the amount by 700,000

OpenStudy (amistre64):

sounds like an economics question ... i recall there being a factor expression for some of this stuff

OpenStudy (amistre64):

or im reading to much into the question and it simply wants to know the difference between the percentages that have to be kept in reserve

OpenStudy (amistre64):

35k (5-2.5)%

OpenStudy (amistre64):

lol, all the options are bigger than the amount in the question, so either youve got a typo, or my memory is calling up something else

OpenStudy (anonymous):

Its on my test so its not a typo thats just what it says

OpenStudy (amistre64):

well, then my initial thought was right, it has something to do with a factor 35000*(.95) = 33250 to lend 33250*(.95) = 31587.5 to lend 31587.5*(.95) = 30008.125 to lend adding all those up till you get to about zero to lend is the amount of money a 5% restriction stimulates the economy

OpenStudy (amistre64):

does that ring a bell? or am i off again?

OpenStudy (dumbcow):

i think its simpler than that, 35k is the reserve amount the bank cannot lend out if rate is 5% , then 35000 = .05X X = 700,000 if rate is 2.5%, then 35000 = .025X X = 1,400,000 thus amount increases by 700,000

OpenStudy (amistre64):

35k is the amount to be reserved eh ..... that at least gets us in the option ballpark :) i still cant inerpret that from the given :/

OpenStudy (amistre64):

ironically, my idea gets me about the same

OpenStudy (dumbcow):

it is a poorly worded question i think .... it was nice to have the answer options available :)

OpenStudy (amistre64):

\[A_n=35000(.95)^{n-1}\] assuming 1 dollar is a good stopping point \[1=35000(.95)^{n-1}~;~n=205\] \[1=35000(.975)^{n-1}~;~n=415\] \[35000(\frac{1-.975^{415}}{1-.975}-\frac{1-.95^{205}}{1-.95})=699,980.70\]

OpenStudy (amistre64):

so an increase of about 700,000 would be good :)

OpenStudy (dumbcow):

nice :), however present value equations probably isn't in scope of the class

OpenStudy (amistre64):

lol, maybe it should be ;)

OpenStudy (dumbcow):

to get exact numbers you need n-> infinity resulting in an infinite geometric sum yielding \[35,000(\frac{1}{1-.975} - \frac{1}{1-.95}) = 35000(40-20) = 700,000\] ha